Find the values of x and y in parallelogram ABCD.
AB= 2y, BC= y + 3, CD= 5x-1, DA= 2x+4

Answer by stanbon(75887)

Find the values of x and y in parallelogram ABCD.
AB= 2y, BC= y + 3, CD= 5x-1, DA= 2x+4
--------
Equations:
AB = CD
AD = BC
---------------
2y = 5x-1
2x+4 = y+3
--------
Rearrange:
5x - 2y = 1
2x - y = -1
------------------
Modify for elimination::
5x - 2y = 1
4x - 2y = -2
--------------------
Subtract and solve for "x":
x = 3
----
Solve for "y":
2y = 5x-1
2y = 15-1
y = 7
---------------------
AB= 2y = 14
BC= y + 3 = 10
CD= 5x-1 = 14
DA= 2x+4 = 10
=================
Cheers,
Stan H.
=================